DCPAM Reference Manual

GFD Dennou Club

平成 17 年 2 月 18 日

目次

1 Module axis type mod 2 1.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Error Handling . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 Known Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.4 Note . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.5 Future Plans . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.6 Dependency . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.7 Public Interface . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.8 Generic Procedure . . . . . . . . . . . . . . . . . . . . . . . . 2 1.9 Derived Types . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.10 Procedure Interface . . . . . . . . . . . . . . . . . . . . . . . 3

1.10.1 Copy AXISINFO data (for 0 dimensional data) … 3 1.10.2 Copy AXISINFO data (for 1 dimensional data) … 3 1.10.3 Copy AXISATTR (in gt4f90io) data (for 1 dimen-

sional data) . . . . . . . . . . . . . . . . . . . . . . . . 4 1.10.4 Initialize AXISATTR (in gt4f90io) data (for 0 dimen-

sional data) . . . . . . . . . . . . . . . . . . . . . . . . 4 1.10.5 Initialize AXISATTR (in gt4f90io) data (for 1 dimen-

sional data) . . . . . . . . . . . . . . . . . . . . . . . . 5

1 Module axis type mod

• Developers: Morikawa Yasuhiro

• Version: $Id: axis type.f90,v 1.8 2005/01/19 08:52:24 morikawa Exp $

• Tag Name: $Name: $

• Change History:

1.1 Overview

This module provide derived types including all information about Axes data, and utility to treat the derived type variables.

座標軸に関する全ての情報を包括する構造体、およびその構造体変数を扱うためのユーティリティを提供する。

1.2 Error Handling

1.3 Known Bugs

1.4 Note

1.5 Future Plans

1.6 Dependency

use type_mod,    only: REKIND, DBKIND, INTKIND, TOKEN, STRING
use gt4_history, only: GT_HISTORY_AXIS, GT_HISTORY_ATTR

1.7 Public Interface

private
public :: AXISINFO                        ! derived types
public :: axis_type_copy, axis_attrs_copy ! subroutines
public :: axis_attrs_init                 ! subroutines

1.8 Generic Procedure

interface axis_type_copy
   module procedure axis_type_copy0, axis_type_copy1

end interface
interface axis_attrs_init
   module procedure axis_attrs_init0, axis_attrs_init1
end interface

1.9 Derived Types

次元に関する情報、およびデータを格納

type AXISINFO
   type(GT_HISTORY_AXIS)             :: axisinfo ! gt4 変数情報
   type(GT_HISTORY_ATTR), allocatable:: attrs(:) ! 属性情報群
   real(DBKIND)         , allocatable:: a_Dim(:) ! 次元データ
   logical                :: stored = .false.  ! 格納完了フラグ
end type AXISINFO

1.10 Procedure Interface

1.10.1 Copy AXISINFO data (for 0 dimensional data)

subroutine axis_type_copy0(from, to)

Dependency

use type_mod, only: STRING
use dc_trace, only: BeginSub, EndSub

Input

type(AXISINFO), intent(in)  :: from

Output

type(AXISINFO), intent(out) :: to

1.10.2 Copy AXISINFO data (for 1 dimensional data)

subroutine axis_type_copy1(from, to)

Dependency

use type_mod, only: STRING, INTKIND
use dc_trace, only: BeginSub, EndSub

Input

type(AXISINFO), intent(in)  :: from(:)

Output

type(AXISINFO), intent(out) :: to(:)

1.10.3 Copy AXISATTR (in gt4f90io) data (for 1 dimensional

data)

subroutine axis_attrs_copy(from, to)

Dependency

use type_mod, only: STRING, INTKIND
use gt4_history, only: GT_HISTORY_ATTR
use dc_trace, only: BeginSub, EndSub, DbgMessage

Input

type(GT_HISTORY_ATTR), intent(in)  :: from(:)

Output

type(GT_HISTORY_ATTR), intent(out) :: to(:)

1.10.4 Initialize AXISATTR (in gt4f90io) data (for 0 dimen-

sional data)

subroutine axis_attrs_init0(attrs)

Dependency

use type_mod, only: STRING, INTKIND
use gt4_history, only: GT_HISTORY_ATTR
use dc_trace, only: BeginSub, EndSub, DbgMessage

In/Out

type(GT_HISTORY_ATTR), intent(inout):: attrs

1.10.5 Initialize AXISATTR (in gt4f90io) data (for 1 dimen-

sional data)

subroutine axis_attrs_init1(attrs)

Dependency

use type_mod, only: STRING, INTKIND
use gt4_history, only: GT_HISTORY_ATTR
use dc_trace, only: BeginSub, EndSub, DbgMessage

In/Out

type(GT_HISTORY_ATTR), intent(inout):: attrs(:)